OpenStudy (anonymous):

What is 1 3/14 - (-2 1/7)?

OpenStudy (caozeyuan):

well, -- is +, right?

OpenStudy (anonymous):

1 3/14 minus (-2 1/7)

OpenStudy (refrac532):

That can be simplified to...?

OpenStudy (anonymous):

I got -1 1/13 for my answer, correct?

OpenStudy (refrac532):

Do you mean -1 1/14?

OpenStudy (anonymous):

uh no

OpenStudy (refrac532):

Then what are you doing xD

OpenStudy (caozeyuan):

you have two - thats just a +

OpenStudy (anonymous):

Ok I'll show you what I did

OpenStudy (refrac532):

You can change 1/7 to match the other denominator

OpenStudy (caozeyuan):

if you minus a number you just add the opposite, so if you minus negative you add positive and thats easier

OpenStudy (refrac532):

For example, 1 - (-3) can be changed to 1 + 3, or 4.

OpenStudy (anonymous):

1 3/14 = 15/7 (-2 1/7) = 17/14 LCD Is 14 17/14 - 30/14= -13/14 Divide and get -1 1/13

OpenStudy (caozeyuan):

so many error, what are you doing?!

OpenStudy (refrac532):

No, you need to add the mixed fractions.

OpenStudy (refrac532):

Are you sure you know what you're doing?

OpenStudy (caozeyuan):

add them up!

OpenStudy (anonymous):

I know what I'm doing i'm supposed to subtract

OpenStudy (caozeyuan):

1+2=3,3/14+2/14=5/14, so answer is 3 5/14

OpenStudy (anonymous):

isn't that correct???

OpenStudy (caozeyuan):

ok, basics, what is 1-(-1)?

OpenStudy (anonymous):

Oh nvm I get it now

OpenStudy (anonymous):

I forgot about the subtracting rule thing

OpenStudy (refrac532):

I think you also forgot about common denominators

OpenStudy (refrac532):

In no way is 13 related to 7 and 14

OpenStudy (caozeyuan):

np, I brain fart a lot too

OpenStudy (anonymous):

@caozeyuan So is 3 5/14

OpenStudy (caozeyuan):

great!